

********************************************************************
UARS WEEKLY STATUS REPORT - 12 MAR 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar, offset, and occultation
data without apparent problems.
Reprocessing of the SUSIM data to data version 20 continues.
On March 12, 1999, SUSIM begins a major calibration using all four
deuterium calibration lamps. Afterward, SUSIM will begin a
fundamentally new operational scenario in which elements of the optical
path used for daily observations are replaced in favor of ones which are
more responsive. It is expected that this change will increase the
quality of the received signals while maintaining SUSIM's long-term
calibration.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
The instrument is now off as scheduled. MLS is scheduled to begin north viewing
operations on Apr 4 and operate though Apr 10.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing recived.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
49 days PEM (3.32) PROD (OA-L2) (760R-229C=531L)
145 days SUS (7.0) PROD (668R-401C=267L)
11 days WIN (5.11) PROD TEST (LOG) (11R-11C=0L)
11 days WIN (5.11) PROD TEST (L1-L3) (11R-11C=0L)
Preventive Maintenance for the CDHF systems and peripherals is due in March.
It is scheduled for Tuesday 03/16/99.
The CDHF is scheduled to participate in an End to End Y2K test on April
5th. This test will require a system clock time shift and an interruption to
normal operations.
CDHF pending software upgrades: DEC AXP Fortran 6.5, DEC AXP OPEN
VMS 6.2, OpenIngres
SOLSTICE Wall Clock time increase/no change in status.
***********
FOT
***********
REPORTING PERIOD
This report covers 27 February 1999 (Orbit 40781, GMT Day 058 through 05 March.
Observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
MLS 058/0037 40781 On
SOLSTICE 058/0052 40781 Day Only
MLS 063/2320 40870 Off
ACRIM 064/0100 40870 On
SOLSTICE 064/0112 40870 Full Time
WINDII 064/0432 40873 On
SUSIM 064/1205 40878 On
The beta angle ranged between the angles of 22.9 degrees to 46.2 degrees
for this report period. The beta angle is now increasing toward a maximum
angle of +53 degrees on 10 March 1999 (DOY 069).
A +0.5 Degree Roll Offset Maneuver was performed on 04 March (DOY 063) in
Orbit 40870 at 063/22:48:00. The maneuver was requested by the WINDII
instrument to facilitate the observation of Polar Mesospheric Clouds. UARS
will return to level flight at the end of day on 16 March (DOY 075). The
exact time for the return maneuver has not yet been finalized.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 1.0 to 1.3 degrees C. Additional battery performance data
for each battery on 26 February 1999 and 05 March 1999 are provided in
Appendix D.
The clock error ranged from 10.7 to -12.6 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 02
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 03
% MLS 05
% WINDII 01
% ACRIM 01
% HALOE 01
% SUSIM 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 06
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 99/054 - 2287.496948
Transponder B: 99/055 - 2287.498731
GROUND SYSTEM OPERATION
Six (6) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 99/063 show the observatory data loss to be
22 hours, 46 minutes, 32.756 seconds (an increase of 05.152 seconds since
last report period). This is a 0.0351 percent data loss which equals a
99.9649 percent data capture for the mission.
The increase of 05.152 seconds occurred over five (5) days:
7 1.120 seconds on Day 058 due to no overlap, data hit.
7 0.224 seconds on Day 059 due to data hit.
7 1.152 seconds on Day 061 due to no overlap.
7 0.672 seconds on Day 062 due to data hit.
7 1.984 seconds on Day 063 due to data hit (IP).
No additional Y2K "test" 8-week planning aids were received by the FOT this
week. The FOT is still awaiting the delivery of two "corrected" planning
aids to complete the test set. These two remaining files will most likely
be delivered to the FOT next week, after which time the FOT will process
the files and make them available to the MPG and Instrument Operations
Teams.
The UARS Test and Training Simulator (UTTS) remains non-operational. The
estimated time to repair the UTTS and return it to operation is unknown.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continues to support the PEM team with diagnosis of the Energy
Processor 6 status error. PEM EP6 is still not coming up nominally under
PMON control. The FOT will conduct additional tests to identify where in
the turn-on sequence the conflict occurs in the coming week.
The FOT continues to investigate the feasibility of a new PMON action to
change battery V/T curves based on loading to mitigate battery risks in sun
pointing safehold.
YEAR 2000 OPERABILITY SUPPORT - No change
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
Planning continues for the simulator testing.
UARS RE-ENGINEERING - TELEMETRY PROCESSING - No change
MEETINGS
FOT participated in a meeting on 03 March 1999 concerning the development
of an Operational Concept Document for the proposed Integrated Mission
Operation Complex (IMOC).
FOT attended a Y2K Contingencies Meeting hosted by John Valiant on 03 March
1999. The FOT turned in a questionnaire about possible UARS Y2K scenarios
and received a list of the priority order for spacecraft should a crisis
arise.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver (forward-to-reverse) is scheduled for 28
March 1999 (DOY 087).
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 05 March 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
40781 058/0037 99-080 MLS Power ON
40796 059/0037 99-081 OBC Clk Adj fr 0.00 to 20.6 msec/day
40797 059/0223 99-082 MLS MLS Stop/Run
40823 060/1952 99-083 OBC Clk Adj fr 20.6 to 0.00 msec/day
40823 060/1952 99-084 MLS Scan actuator off due to HI temps.
40825 060/2132 99-083 MLS Restart Scan sequence
40858 063/0428 99-086 OBC Clk Adj fr 0.00 to 20.6 msec/day
40870 063/2320 99-087 MLS Power OFF
40870 063/2320 99-088 HALOE Update Sunrise parameter
40871 064/0100 99-089 ACRIM Power ON
40873 064/0432 99-090 WINDII Power ON
40878 064/1205 99-091 SUSIM Power ON
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 05 March 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
NBRECBI Y-HI NBTR B Current 40786
QTSSPPPX Y-HI SSPP at +x therm 40833,61,73
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 05 March 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
40796 059/00:43 (27 Feb 99) from 00.0 to 20.6 msec/day
40823 060/19:56 (01 Mar 99) from 20.6 to 00.0 msec/day
40858 063/04:33 (03 Mar 99) from 00.0 to 20.6 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 05 March 1999
26 Feb 1999 - GMT Day 057 Beta = 22.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -5.6/-33.6 -3.05/-4.31 +0.0/ +0.0 16.8 0.000 0.0
2 +61.6/ 0.0 +5.47/+4.51 +31.6/-14.0 25.4 1.024 21.2
3 +67.2/-16.8 +4.51/+3.23 +31.2/-15.2 25.4 1.024 21.0
05 Mar 1999 - GMT Day 064 Beta = 46.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -11.2/-33.6 -3.05/-4.62 +0.0/ +0.0 16.8 0.000 0.0
2 +22.4/-16.8 +5.15/+3.87 +28.0/-15.2 26.2 1.025 26.0
3 +39.2/-11.2 +3.87/+2.92 +27.2/-16.0 26.1 1.021 25.7
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 05 March 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-006 FOT 40802 059/1012 14170 Generic Late Acquisition
99-006 FOT 40825 060/2132 14170 Generic Late Acquisition
99-006 FOT 40836 061/1655 14170 Generic Late Acquisition
99-006 FOT 40851 062/1701 14170 Generic Late Acquisition
99-006 FOT 40855 062/2317 14170 Generic Late Acquisition
99-006 FOT 40865 063/1644 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-006 FOT see list Mar 99 14170 Generic Late Acq
SEVNTFW99/10